please write the answer like this
answer=?
Betty DeRose, Inc. operates two departments, the handling department and
the packaging department. During April, the handling department reported
the following information:

                                           % complete      % complete
                                units         DM           conversion
work in process, April 1        18,000        38%             71%
units started during April      80,000
work in process, April 30       44,000        28%             74%

The cost of beginning work in process and the costs added during April
were as follows:

                                 DM         Conversion       Total cost
work in process, April 1      $ 64,996       $152,785         $217,781
costs added during April       190,336        232,407          422,743
total costs                    255,332        385,192          640,524

Calculate the cost of goods completed and transferred out during April
using the weighted average process costing method.
Answer #1

Unit transferred out = 18000+80000-44000 = 54000

DM Conversion
Unit transferred out 54000 54000
Ending WIP 44000*28% = 12320 44000*74% = 32560
Equivalent unit 66320 86560

Unit cost

DM Conversion
Beginning WIP 64996 152785
Cost added 190336 232407
Total Cost 255332 385192
Equivalent unit 66320 86560
Unit cost 3.85 4.45

Cost of unit transferred out

DM Conversion Total
Cost of unit transferred out 54000*3.85 = 207900 240300 448200